OpenStudy (anonymous):

4x + 6y = 14 -2x + 2y = -2

OpenStudy (anonymous):

i really need help! im doing credit recovery and i dont know how to do any of this. Please help!!

OpenStudy (anonymous):

okay so

OpenStudy (anonymous):

we have to get all the x's and all the y's on the same side, so we have: 4x + 6y = 14 -2x + 2y = -2 4x - 2x = 2x, and 6y + 2y = 8y. and 14 - 2 is 12. so we get 2x + 8y = 12

OpenStudy (anonymous):

so basically what i did was combine the terms

OpenStudy (anonymous):

i'm not sure if we're supposed to solve or simplify but that is it simplified don't know how to solve lol but i hope that helped some